OpenStudy (anonymous):

solve: m/(2m+2)=-2m/(4m+4)+2m-3/(m+1)

OpenStudy (anonymous):

-2m/(4m+4)=-2m/4(m+1)=-0.5m/(m+1) then ;r.h.s =(-0.5m+2m -3)/(m+1) ten r.h.s =(1.5m-3)/(m+1) at L.h.s =m/2(m+1) by multiply by (m+1) at both sides then 0.5m=1.5m-3 so M=3